# An Efficient Wireless iBCI Headstage with Adaptive ADC Sample Rate

### 23 Apr 2026 v1

@@ I Introduction
Implantable Brain-Computer Interfaces (iBCIs) have emerged as a transformative technology for restoring functional independence in individuals with severe motor impairments, such as tetraplegia and amyotrophic lateral sclerosis (ALS) [18]. Fig. 1 displays the basic workflow of typical iBCI systems. By employing high-density microelectrode arrays (MEAs) [14, 19] to record extracellular action potentials, iBCIs bypass damaged neuromuscular pathways to establish a direct communication link between the brain and external prosthetic devices. Recent clinical breakthroughs [6] have demonstrated the remarkable capability of iBCIs to decode complex motor intent with high degrees of freedom, enabling users to control robotic limbs [24], communicate with brain-to-text decoders [23], and even regain some degree of tactile feedback [3]. As the field transitions from proof-of-concept laboratory demonstrations to long-term clinical applications [5], the pursuit of high-fidelity, large-scale neural recording remains the cornerstone of sophisticated neural decoding.
Despite iBCIs’ potential, the clinical translation of current iBCI systems is severely impeded by their reliance on percutaneous tethers. These physical connections not only introduce a persistent risk of infection at the skin interface but also anchor users to laboratory hardware, restricting mobility and natural behavior. Consequently, there is a critical imperative to transition toward fully implantable, wireless systems capable of supporting long-term home use. However, untethering the device introduces a formidable ”bandwidth bottleneck.” As recording densities scale to thousands of channels to enhance decoding fidelity, the aggregate raw data throughput rapidly outstrips the capacity of energy-efficient telemetry. While emerging wireless technologies such as Terahertz (THz) [9, 13, 12] and Ultra-Wideband (UWB) [8] theoretically support gigabit-level data rates, their practical utility is strictly capped by the power dissipation limits necessary to prevent thermal damage to the surrounding cortical tissue.
@@
To mitigate the bandwidth bottleneck in high-density iBCIs, extensive research has focused on reducing data throughput via on-chip or server-assisted compression methodologies. Conventional strategies typically rely on digital-side processing, such as energy-based spike detection [16], Compressive Sensing (CS) leveraging transform-domain sparsity [26], or dimensionality reduction techniques like Principal Component Analysis (PCA) and autoencoder-based feature extraction [25]. While effective at reducing transmission rates, these methods often impose a paradox: effective compression requires sophisticated on-chip learning of signal characteristics, yet resource-constrained headstages lack the power and memory limits to support such computation without compromising feature extraction accuracy or signal fidelity. Recent server-driven architectures [11, 10] attempt to alleviate this by offloading the heavy learning burden to external processors to determine electrode-specific compression levels; however, these schemes operate largely at the application level and rely on on-chip process after the full-resolution ADC sampling. This ”digitize-then-compress” paradigm is inherently inefficient, as it expends energy digitizing clinically irrelevant noise before discarding it. Consequently, there is a critical need for approaches that shift optimization to the hardware front-end, regulating data volume directly at the acquisition source (e.g., the ADC) to maximize energy efficiency while preserving essential neural information.
@@
  • We design a lightweight server-side predictor that estimates spike detection error from spike templates under candidate sampling rates and thresholds, enabling aggressive rate reduction while bounding accuracy loss.
@@ II-B Spike Sorting
Spike sorting serves as a foundational computational process in neurophysiology, designed to identify the firing events from each neuron; however, this procedure entails substantial computational complexity, particularly when applied to large-scale datasets. Traditional spike sorting algorithms can be categorized into feature extraction [22] and template matching [15]. As high-density MEAs [19] containing hundreds or thousands of channels, the analysis faces the critical challenge of signal superposition. In these dense recording environments, the probability of ”spatiotemporal collisions”, where waveforms from adjacent neurons overlap and interfere increases significantly, rendering traditional clustering methods inadequate. Consequently, the field has shifted towards template-matching algorithms, most notably exemplified by frameworks such as Kilosort [15]. These approaches treat the raw data as a linear sum of individual neuronal signatures to resolve overlapping events efficiently. In this context, a ”template” is defined as the high-dimensional summary of a specific neuron’s electrical footprint; explicitly, it represents the average waveform profile of that neuron distributed across all relevant recording channels, capturing its unique spatiotemporal signature within the array. Fig. 2 display four spike templates from a rat 384-electrode Neuropixel dataset using Kilosort.
@@ II-C Neural Decoder
Alternatively, recent studies have demonstrated the efficacy of extracting high-frequency band power (e.g., 300-1000 Hz) as a surrogate for precise spiking activity (Even-Chen et al., 2020; Nason et al., 2020). While this feature-engineering approach permits aggressively low sampling rates (e.g., 2 kHz) and significantly reduces data bandwidth, our work focuses on preserving the precise waveform and spike timing of single- or multi-unit activities. This fine temporal resolution is essential for high-precision temporal decoding and advanced template matching, distinguishing our adaptive sampling strategy from band-power extraction methods.
@@ III-A Overview
Fig. 3 shows the workflow of the proposed framework, which operates in two phases: offline calibration and online streaming. During offline calibration, the system first extracts spike templates to characterize the neural activity observed on each electrode. These templates are sent to the server, where a lightweight predictor estimates the spike detection loss associated with a candidate electrode configuration $(s_{i},th_{i})$, where $s_{i}$ and $th_{i}$ denote the sample rate and detection threshold of electrode $i$, respectively. Based on the predicted loss, the server selects an electrode-specific configuration that minimizes the sampling rate while satisfying a user-defined error budget.
@@ III-B Motivation
Why adaptive sample rates and thresholds are feasible. We begin by examining spikes recorded from a single electrode. We observe that spikes with larger amplitudes remain detectable even when the sampling rate is reduced, whereas weaker spikes are more sensitive to downsampling. Fig. 4 shows two representative examples under the same detection threshold. Spike A remains detectable with a downsampling factor of four, while Spike B, which has a smaller amplitude, can tolerate a factor of only three before its threshold crossing is lost. Because each electrode records a different mixture of neurons and noise, the minimum sampling rate required to preserve spike detection varies substantially across electrodes. This electrode-to-electrode variability creates the opportunity for adaptive optimization.
@@ III-C1 Problem Formulation
    False negative rate (FNR): the fraction of true spikes that are missed because the sampling rate is too low or the threshold is too strict.
@@ III-C2 Lightweight Predictor
Input and output. The input is the concatenation of the candidate sample rate $s_{i}$, the detection threshold $th_{i}$, and the template waveform $\mathcal{P}_{i}$. The network outputs two scalars: the predicted false negative rate $\hat{e}_{FNR}$ and the predicted false positive rate $\hat{e}_{FPR}$.
Architecture. The predictor is a compact multilayer perceptron with two hidden layers of 48 and 24 neurons, respectively, followed by ReLU activations. This design is intentionally lightweight: it is expressive enough to capture the nonlinear relationship between waveform shape, threshold, sampling rate, and detection error, while remaining fast enough for practical server-side optimization.
@@ III-C3 Training and Optimization
The value of $\epsilon$ depends on the robustness of the downstream decoder. For conventional decoders, we use $\epsilon=5\%$. For semantic decoders such as Brain-to-Text systems that rely on large language models (LLMs) to absorb moderate front-end errors, the constraint can be relaxed to $\epsilon=10\%$ to enable more aggressive compression and larger power savings.
Figure 5: Prediction error of the spike missing rate (left) and spike mis-detection rate (right) on rat and human datasets.
We evaluate the predictor on a rat dataset [7] and a human cortex dataset [21], both recorded with Neuropixels MEAs [19]. Fig. 5 shows the prediction error distribution across electrodes. The predictor achieves errors below 2% for the spike missing rate and below 0.6% for the spike mis-detection rate. Importantly, this model is trained only once offline and runs on the server, so it introduces no additional area or power overhead to the headstage. The inference latency is 4.3 ms, and the optimization of $\{s_{i},th_{i}\}$ converges within 46 s on an Intel Core i7-13500KF CPU, which is sufficiently fast to track template drift that typically occurs over tens of minutes [11].
@@ III-D ADC-Level Sample Rate Adjustment
Let $\mathbf{S}=\{s_{1},s_{2},\dots,s_{N}\}$ denote the target sample-rate vector, where $s_{i}$ is the desired sampling rate of electrode $i$. The ADC interface is driven by a master clock $f_{clk}$ and executes a periodic command schedule. If each execution round contains $N_{t}$ clock cycles, among which $N_{s}$ cycles correspond to active sampling commands, then the maximum achievable sampling rate is
@@
To realize a lower sampling rate, each electrode is assigned an integer downsampling factor $x_{d,i}$ from a supported set of hardware factors. The FPGA then samples electrode $i$ once every $x_{d,i}$ scheduling rounds. The implemented rate is therefore
@@ IV Implementation
Electrode-Level Ground Truth Construction. Our adaptive ADC controller operates at the electrode level, whereas spike sorting provides annotations at the neuron level. More specifically, spike sorting yields source-resolved labels that indicate whether neuron $n\in\mathcal{N}$ fires at time $t$, while our ADC rate adaptation requires sensor-resolved labels that indicate whether electrode $e$ should observe a detectable spike event at time $t$. To bridge this granularity mismatch, we project neuron-level spike labels onto the electrode array.
Let $y_{n}(t)\in\{0,1\}$ denote the binary spike train of neuron $n$, and let $\mathcal{E}_{n}$ denote its spatial footprint, i.e., the set of electrodes on which the waveform of neuron $n$ is visible above the noise floor. We define the electrode-level ground truth $G_{e}(t)$ as
$G_{e}(t)=\mathbb{I}\!\left(\sum_{n\in\mathcal{N}}y_{n}(t)\,\mathbb{I}(e\in\mathcal{E}_{n})>0\right)=\bigvee_{n:e\in\mathcal{E}_{n}}y_{n}(t),$ (8)
where $\mathbb{I}(\cdot)$ is the indicator function. Therefore, $G_{e}(t)=1$ whenever at least one neuron whose activity is visible on electrode $e$ fires at time $t$. If multiple neurons fire simultaneously and are observable on the same electrode, they are still counted as a single positive detection event. This formulation matches the behavior of our per-electrode threshold detector and ensures that the optimization target is consistent with the control granularity of the ADC scheduler.
Figure 6: Headstage prototype.
Headstage Architecture and Runtime Data Flow. Fig. 6 shows our wireless headstage prototype, whose total power consumption is below 173 mW. The system consists of three cascaded modules: an acquisition front-end (Intan RHD2132), a processing and control core (Digilent Cmod A7 FPGA), and a wireless transceiver (ESP32-S3). These modules operate in a server-guided closed loop.
At the beginning of each adaptation interval, the server computes the electrode-wise configuration vector
$\mathbf{C}=\{(s_{i},th_{i})\}_{i=1}^{N},$ (9)
where $s_{i}$ and $th_{i}$ denote the target sample rate and detection threshold for electrode $i$, respectively. The server then transmits $\mathbf{C}$ to the headstage through the wireless downlink. Upon reception, the FPGA translates $\mathbf{C}$ into a concrete per-electrode sampling schedule and detector configuration.
Acquisition: The FPGA drives the Intan RHD2132 through SPI and selectively issues acquisition commands according to the configured schedule. Electrodes assigned lower sample rates are sampled only in their scheduled rounds, such that rate reduction is enforced directly at the acquisition stage rather than after digitization. The resulting digitized samples are streamed from the front-end to the FPGA in real time.
On-chip processing: The FPGA serves as the main control and computation layer. For each acquired stream, it executes a real-time signal processing pipeline consisting of band-pass filtering, whitening, and threshold-based spike detection. Since downsampling is already applied during acquisition, the FPGA only processes samples that are physically collected, thereby avoiding the overhead of post-digitization decimation on discarded data.
Transmission: Instead of transmitting raw broadband neural signals, the FPGA packetizes only the detected spike events and transfers these event packets to the ESP32-S3 over SPI. The ESP32-S3 then wirelessly transmits the compressed event stream to the server for downstream decoding.
Overall, this implementation keeps the headstage lightweight: the computationally intensive optimization is performed on the server, while the on-device hardware is responsible only for schedule execution, real-time signal conditioning, spike detection, and low-bandwidth event transmission.
@@ V Evaluation
V-A Methodology.
Datasets. We evaluate the prototype of the headstage with five iBCI datasets. Three raw neural recordings across macaque [2], rat [7] and human [21] with Utah and Neuralpixel MEAs to assess the spike detection performance. Besides, we also use a visual and a motor decoders from CEBRA [17] to assess the impact of downsampling on the downstream tasks.
Metrics. We propose to use compression ratio and spike detection error to evaluate the efficiency of adaptive sampling system. Besides, we measure the power consumption and resource utilization on FPGA to profile the overhead.
Baselines We benchmark our approach against two neural signal compression algorithms implemented on the FPGA: Discrete Cosine Transform (DCT) [1] and Compressive Sensing (CS) [26]. Both schemes represent application-layer implementations that process fully digitized neural signals.
Evaluation Metrics We assess system performance across two dimensions: (1) Signal Fidelity: We quantify the Compression Ratio (CR) and the Spike Detection Error (SDE) (comprising false negatives and false positives) to measure data efficiency and integrity. (2) Hardware Efficiency: We profile the power consumption (mW) and FPGA resource utilization (LUTs/Flip-Flops) to determine the implementation overhead.
Experimental Setup We conduct a comprehensive data-driven evaluation. Datasets recorded with 100- and 384-electrode MEAs were partitioned into 32-channel subsets to match the acquisition capacity of the Intan system.
V-A1 ADC Simulation
@@
V-A2 Stochastic Error Injection for Decoders
For the CEBRA decoding benchmarks, raw neural traces are unavailable, precluding direct resampling. Instead, we assess robustness by injecting the empirical errors observed in the Rat dataset. We perturb the ground truth spike sequences $\mathbf{y}$ to generate corrupted traces $\hat{\mathbf{y}}$ as follows: (1) Missed Spikes: True spikes are removed (set to 0) with a probability equal to the empirical missed spike rate of the target electrode observed in the rat dataset. (2) Mis-detection spikes: Spurious spikes are inserted (0 flipped to 1) at random idle timesteps with a probability derived from the empirical mis-detection rate of the corresponding electrode in the rat dataset.

### 24 Apr 2026 v2

@@ I Introduction
Implantable Brain-Computer Interfaces (iBCIs) have emerged as a transformative technology for restoring functional independence in individuals with severe motor impairments, such as tetraplegia and amyotrophic lateral sclerosis (ALS) [21, 18]. Fig. 1 displays the basic workflow of typical iBCI systems. By employing high-density microelectrode arrays (MEAs) [16, 22] to record extracellular action potentials, iBCIs bypass damaged neuromuscular pathways to establish a direct communication link between the brain and external prosthetic devices. Recent clinical breakthroughs [6] have demonstrated the remarkable capability of iBCIs to decode complex motor intent with high degrees of freedom, enabling users to control robotic limbs [27], communicate with brain-to-text decoders [26], and even regain some degree of tactile feedback [3]. As the field transitions from proof-of-concept laboratory demonstrations to long-term clinical applications [5], the pursuit of high-fidelity, large-scale neural recording remains the cornerstone of sophisticated neural decoding.
Despite iBCIs’ potential, the clinical translation of current iBCI systems is severely impeded by their reliance on percutaneous tethers. These physical connections not only introduce a persistent risk of infection at the skin interface but also anchor users to laboratory hardware, restricting mobility and natural behavior. Consequently, there is a critical imperative to transition toward fully implantable, wireless systems capable of supporting long-term home use. However, untethering the device introduces a formidable ”bandwidth bottleneck.” As recording densities scale to thousands of channels to enhance decoding fidelity, the aggregate raw data throughput rapidly outstrips the capacity of energy-efficient telemetry. While emerging wireless technologies such as Terahertz (THz) [11, 15, 14, 10] and Ultra-Wideband (UWB) [8, 9] theoretically support gigabit-level data rates, their practical utility is strictly capped by the power dissipation limits necessary to prevent thermal damage to the surrounding cortical tissue.
@@
To mitigate the bandwidth bottleneck in high-density iBCIs, extensive research has focused on reducing data throughput via on-chip or server-assisted compression methodologies. Conventional strategies typically rely on digital-side processing, such as energy-based spike detection [19], Compressive Sensing (CS) leveraging transform-domain sparsity [29], or dimensionality reduction techniques like Principal Component Analysis (PCA) and autoencoder-based feature extraction [28]. While effective at reducing transmission rates, these methods often impose a paradox: effective compression requires sophisticated on-chip learning of signal characteristics, yet resource-constrained headstages lack the power and memory to support such computation without compromising feature extraction accuracy or signal fidelity. Recent server-driven architectures [13, 12] attempt to alleviate this by offloading the heavy learning burden to external processors to determine electrode-specific compression levels; however, these schemes operate largely at the application level and rely on on-chip processing after the full-resolution ADC sampling. This ”digitize-then-compress” paradigm is inherently inefficient, as it expends energy digitizing clinically irrelevant noise before discarding it. Consequently, there is a critical need for approaches that shift optimization to the hardware front-end, regulating data volume directly at the acquisition source (e.g., the ADC) to maximize energy efficiency while preserving essential neural information.
@@
  • We design a lightweight server-side predictor that estimates spike detection error from spike templates under candidate sample rates and thresholds, enabling aggressive rate reduction while bounding accuracy loss.
@@ II-B Spike Sorting
Spike sorting serves as a foundational computational process in neurophysiology, designed to identify the firing events from each neuron; however, this procedure entails substantial computational complexity, particularly when applied to large-scale datasets. Traditional spike sorting algorithms can be categorized into feature extraction [25] and template matching [17]. With high-density MEAs [22] containing hundreds or thousands of channels, the analysis faces the critical challenge of signal superposition. In these dense recording environments, the probability of ”spatiotemporal collisions”, where waveforms from adjacent neurons overlap and interfere increases significantly, rendering traditional clustering methods inadequate. Consequently, the field has shifted towards template-matching algorithms, most notably exemplified by frameworks such as Kilosort [17]. These approaches treat the raw data as a linear sum of individual neuronal signatures to resolve overlapping events efficiently. In this context, a ”template” is defined as the high-dimensional summary of a specific neuron’s electrical footprint; explicitly, it represents the average waveform profile of that neuron distributed across all relevant recording channels, capturing its unique spatiotemporal signature within the array. Fig. 2 displays four spike templates from a rat 384-electrode Neuropixels dataset using Kilosort.
@@ III-A Overview
Fig. 3 shows the workflow of the proposed framework, which operates in two phases: offline calibration and online streaming. During offline calibration, the system first extracts spike templates to characterize the neural activity observed on each electrode. These templates are sent to the server, where a lightweight predictor estimates the spike detection loss associated with a candidate electrode configuration $(s_{i},th_{i})$, where $s_{i}$ and $th_{i}$ denote the sample rate and detection threshold of electrode $i$, respectively. Based on the predicted loss, the server selects an electrode-specific configuration that minimizes the sample rate while satisfying a user-defined error budget.
@@ III-B Motivation
Why adaptive sample rates and thresholds are feasible. We begin by examining spikes recorded from a single electrode. We observe that spikes with larger amplitudes remain detectable even when the sample rate is reduced, whereas weaker spikes are more sensitive to downsampling. Fig. 4 shows two representative examples under the same detection threshold. Spike A remains detectable with a downsampling factor of four, while Spike B, which has a smaller amplitude, can tolerate a factor of only three before its threshold crossing is lost. Because each electrode records a different mixture of neurons and noise, the minimum sample rate required to preserve spike detection varies substantially across electrodes. This electrode-to-electrode variability creates the opportunity for adaptive optimization.
@@ III-C1 Problem Formulation
    False negative rate (FNR): the fraction of true spikes that are missed because the sample rate is too low or the threshold is too strict.
@@ III-C2 Lightweight Predictor
Input and output. The input is the concatenation of the candidate sample rate $s_{i}$, the detection threshold $th_{i}$, and the template waveform $\mathcal{P}_{i}$. The network outputs two scalars: the predicted FNR $\hat{e}_{FNR}$ and the predicted FPR $\hat{e}_{FPR}$.
Architecture. The predictor is a compact multilayer perceptron with two hidden layers of 48 and 24 neurons, respectively, followed by ReLU activations. This design is intentionally lightweight: it is expressive enough to capture the nonlinear relationship between waveform shape, threshold, sample rate, and detection error, while remaining fast enough for practical server-side optimization.
@@ III-C3 Training and Optimization
The value of $\epsilon$ depends on the robustness of the downstream decoder. For conventional decoders, we use $\epsilon=5\%$. For more robust decoders that can tolerate moderate front-end errors, the constraint can be relaxed to $\epsilon=10\%$ to enable more aggressive compression and larger power savings.
Figure 5: Prediction error of FNR (left) and FPR (right) on rat and human datasets.
We evaluate the predictor on a rat dataset [7] and a human cortex dataset [24], both recorded with Neuropixels MEAs [22]. Fig. 5 shows the prediction error distribution across electrodes. The predictor achieves errors below 2% for FNR and below 0.6% for FPR. Importantly, this model is trained only once offline and runs on the server, so it introduces no additional area or power overhead to the headstage. The inference latency is 4.3 ms, and the optimization of $\{s_{i},th_{i}\}$ converges within 46 s on an Intel Core i7-13500KF CPU, which is sufficiently fast to track template drift that typically occurs over tens of minutes [13].
@@ III-D ADC-Level Sample Rate Adjustment
Let $\mathbf{S}=\{s_{1},s_{2},\dots,s_{N}\}$ denote the target sample-rate vector, where $s_{i}$ is the desired sample rate of electrode $i$. The ADC interface is driven by a master clock $f_{clk}$ and executes a periodic command schedule. If each execution round contains $N_{t}$ clock cycles, among which $N_{s}$ cycles correspond to active sampling commands, then the maximum achievable sample rate is
@@
To realize a lower sample rate, each electrode is assigned an integer downsampling factor $x_{d,i}$ from a supported set of hardware factors. The FPGA then samples electrode $i$ once every $x_{d,i}$ scheduling rounds. The implemented rate is therefore
@@ IV Implementation
Figure 6: Architecture of the prototype.
Headstage Architecture. Fig. 6 shows our wireless headstage prototype (<<173 mW). It comprises an Intan RHD2132 analog front-end, a Cmod A7 FPGA, and an ESP32-S3 wireless transceiver, operating in a server-guided closed loop.
Each adaptation interval, the server computes and transmits a per-electrode configuration vector $\mathbf{C}=\{(s_{i},th_{i})\}_{i=1}^{N}$, specifying the sample rate $s_{i}$ and detection threshold $th_{i}$ for each electrode. The FPGA drives the Intan front-end via SPI according to the configured schedule, enforcing rate reduction directly at the acquisition stage. Acquired samples pass through an on-chip pipeline of band-pass filtering, whitening, and threshold-based spike detection. Only detected spike events are packetized and wirelessly transmitted to the server via the ESP32-S3, avoiding raw broadband transmission. This design offloads intensive computation to the server while keeping the headstage limited to schedule execution, lightweight signal conditioning, and low-bandwidth event transmission.
@@ V Evaluation
V-A Methodology
Datasets. We evaluate the headstage prototype on five iBCI datasets. Three raw neural recordings from macaque [2], rat [7], and human [24] subjects, acquired with Utah and Neuropixels MEAs, are used to assess spike detection performance. In addition, we use a visual and a motor decoder from CEBRA [20] to assess the impact of downsampling on downstream tasks.
Baselines. We benchmark our approach against two neural signal compression algorithms implemented on the FPGA: Discrete Cosine Transform (DCT) [1] and Compressive Sensing (CS) [29]. Both schemes represent application-layer implementations that process fully digitized neural signals.
Evaluation Metrics. We assess system performance across two dimensions: (1) Signal Fidelity: We quantify the Compression Ratio (CR) and the Spike Detection Error (SDE) (comprising false negatives and false positives) to measure data efficiency and integrity. (2) Hardware Efficiency: We profile the power consumption (mW) and FPGA resource utilization (LUTs/Flip-Flops) to determine the implementation overhead.
Experimental Setup. We conduct a comprehensive data-driven evaluation. Datasets recorded with 100- and 384-electrode MEAs were partitioned into 32-channel subsets to match the acquisition capacity of the Intan system.
V-A1 Electrode-Level Ground Truth Construction
Spike sorting yields neuron-level labels, whereas our ADC controller operates at electrode granularity. To bridge this mismatch, we project neuron-level spike trains onto the electrode array. Let $y_{n}(t)\in\{0,1\}$ denote the spike train of neuron $n$, and let $\mathcal{E}_{n}$ be its spatial footprint, i.e., the set of electrodes on which neuron $n$’s waveform exceeds the noise floor. The electrode-level ground truth is
$G_{e}(t)=\bigvee_{n:\,e\in\mathcal{E}_{n}}y_{n}(t),$ (8)
so that $G_{e}(t)=1$ whenever at least one neuron visible on electrode $e$ fires at time $t$. Simultaneous spikes from multiple neurons on the same electrode are treated as a single detection event, consistent with the per-electrode threshold detector.
V-A2 ADC Simulation
@@
V-A3 Stochastic Error Injection for Decoders
For the CEBRA decoding benchmarks, raw neural traces are unavailable, precluding direct resampling. Instead, we assess robustness by injecting the empirical errors observed in the rat dataset. We perturb the ground truth spike sequences $\mathbf{y}$ to generate corrupted traces $\hat{\mathbf{y}}$ as follows: (1) Missed Spikes (FNR): True spikes are removed (set to 0) with a probability equal to the empirical FNR of the target electrode observed in the rat dataset. (2) Spurious Spikes (FPR): Spurious spikes are inserted (0 flipped to 1) at random idle timesteps with a probability derived from the empirical FPR of the corresponding electrode in the rat dataset.